# Adaptive Querying for Reward Learning from Human Feedback

### v1

@@ IV-A Critical States Selection
IV-B Feedback Format Selection
A feedback format $f^{*}$ that maximizes the information gain is selected to query across $\Omega$. The information gain of a feedback format $f$ at iteration $t$, for $N\!=\!|\Omega|$ critical states, is calculated using the KL divergence between $p^{t}$ and $q^{t}$:
$\mathcal{V}_{f}$ $=\frac{1}{N}\sum_{s\in\Omega}D_{KL}(p^{t}\|q^{t})\cdot\mathbb{I}[f\!=\!f_{H}^{t}]+\mathcal{V}_{f}\cdot(1-\mathbb{I}[f\!=\!f_{H}^{t}]).$ (3)
where, $\mathbb{I}[f\!=\!f_{H}^{t}]$ is an indicator function that checks whether the format in which the human provided feedback, $f_{H}^{t}$, matches the requested format $f$. If no feedback is provided, the information gain of that format remains unchanged.
The following equation is used to select the feedback format $f^{*}$, accounting for feedback cost and user preferences:
$f^{*}=\operatorname*{argmax}_{f\in\mathcal{F}}\underbrace{\frac{\psi(f)}{\mathcal{V}_{f}\cdot\text{C}(f)}+\sqrt{\frac{\log t}{n_{f}+\epsilon}}}_{\Large\text{Feedback utility of $f$}}$ (4)
where $\psi(f)$ is the probability of receiving a feedback in format $f$ and $C(f)$ is the feedback cost, determined using the human preference model $D$. $t$ denotes the current learning iteration, $n_{f}$ is the number of times $f$ was received, and $\epsilon$ is a small value added for numeric stability. Note that $f^{*}$ is the current most informative feedback, based on the formats previously used. This may change as the agent explores and incorporates feedback of other formats. Thus, our feedback selection approach effectively manages the trade-off between selecting feedback formats that were previously used to gather information and examining unexplored formats.
Algorithm 2 outlines our feedback format selection approach. Since the agent has no prior knowledge of NSEs, all state-action pairs in $p$ and $q$ are initialized to safe action label (Line 2). This mapping will be refined based on the feedback received from the human, as learning progresses. It samples $N$ critical states using Algorithm 1 (Line 4). A feedback format $f^{*}$ is selected using Eqn. 4. The agent queries the human for feedback in $f^{*}$ (Line 5). The human provides feedback in format $f^{*}$ with probability $\psi(f^{*})$. Upon receiving the feedback, the agent updates the distribution, $p^{t}$ based on the new NSE information, and trains an NSE prediction model, $\mathcal{P}$ (Lines 6-8). The agent’s belief of NSE distribution $q^{t}$, is updated for all state-action pairs, using $\mathcal{P}$. The information gain $\mathcal{V}_{f^{*}}$ is updated using Eqn. 3, and $n_{f^{*}}$ is incremented (Lines 9-11). This process repeats until the querying budget is exhausted, producing a predictive model of NSEs. Figure 4 illustrates the critical states and the most informative feedback formats selected at each iteration in the vase domain using our approach.
Figure 4: Left: Feedback utility of different formats across iterations. Right: An instance of Vase domain. States marked with a circled number represent the iteration in which the state was identified as a critical state. The color of the circle denotes the feedback format chosen for querying during that iteration.
@@ IV-B Feedback Format Selection
0:  B, Querying budget
0:  $D$, Human preference model
1:  $t\leftarrow 1$; $\mathcal{V}_{f}\leftarrow 0$ and $n_{f}\leftarrow 0,\ \forall f\in\mathcal{F}$
2:  Initialize $p$ and $q$ assuming all actions are safe: $\forall a\in A,\forall s\in S:p(s,a)\leftarrow l_{a}$, $q(s,a)\leftarrow l_{a}$
3:  while $B>0$ do
@@
8:        $\mathcal{P}\leftarrow$ TrainClassifier$(p^{t})$
9:        $q^{t}\leftarrow\{\mathcal{P}(s,a),\forall a\in A,\forall s\in S\}$
@@
(a) Navigation: Unavoidable NSE
(b) Vase: Unavoidable NSE
(c) Safety-gym Push
(d) ATARI Freeway
@@
Figure 5: Top row: average penalty incurred when querying with different feedback selection techniques. Bottom row: illustrations of domains used for evaluation. Red box denotes the agent and the goal location is in green.
V EXPERIMENTS
@@
Baselines (i) Naive Agent: The agent naively executes its primary policy without learning about NSEs, providing an upper bound on the NSE penalty incurred. (ii) Oracle: The agent has complete knowledge about $R_{T}$ and $R_{N}$, providing a lower bound on the NSE penalty incurred. (iii) Reward Inference with $\beta$ Modeling (RI) [33]: The agent selects a feedback format that maximizes information gain according to the human’s inferred rationality $\beta$. (iv) Cost-Sensitive Approach: The agent selects a feedback method with the least cost, according to the preference model $D$. (v) Most-Probable Feedback: The agent selects a feedback format that the human is most likely to provide, based on $D$. (vi) Random Critical States: The agent uses our AFS framework to learn about NSEs, except the critical states are sampled randomly from the entire state space. We use $\theta_{1}\!=\!1$ and $\theta_{2}\!=\!1$ for all our experiments. AFS uses learned $\hat{R}_{N}$. Code will be made public after paper acceptance.
@@
Domains, Metrics and Feedback Formats We evaluate the performance of various techniques on four domains in simulation: outdoor navigation, vase, safety-gym’s push, and Atari freeway. We simulate the feedback for a state-action pair using a softmax action selection [33, 35]: the probability of choosing an action $a^{\prime}$ from a set of all safe actions $A^{*}$ in state $s$ is, $\Pr(a^{\prime}|s)=\frac{e^{Q(s,a^{\prime})}}{\sum\limits_{a\in A^{*}}e^{Q(s,a)}}$. We optimize costs (negations of rewards) and compare techniques using average NSE penalty and average cost to goal, averaged over 100 trials. For navigation, vase and push, we simulate explicit human feedback formats. For Atari we use both explicit (demonstration) and implicit (gaze) feedback from the Atari-HEAD dataset [36]. The penalties for $l_{a}$, $l_{m}$, and $l_{h}$ are 00, $+5$, and $+10$ respectively.
@@
Navigation: In this ROS-based city environment, the robot optimizes the shortest path to the goal location. A state is represented as $\langle x,y,f,p\rangle$, where, $x$ and $y$ are robot coordinates, $f$ is the surface type (concrete or grass), and $p$ indicates the presence of a puddle. The robot can move in all four directions and each costs $+1$. Navigating over grass damages the grass and is a mild NSE. Navigating over grass with puddles is a severe NSE. Features used for training are $\langle f,p\rangle$. Here, NSEs are unavoidable.
@@
Vase: In this domain, the robot must quickly reach the goal, while minimizing breaking a vase as a side effect [4]. A state is represented as $\langle x,y,v,c\rangle$ where, $x$ and $y$ are robot’s coordinates. $v$ indicates the presence of a vase and $c$ indicates if the floor is carpeted. The robot moves in all four directions and each costs $+1$. Actions succeed with probability $0.8$. Breaking a vase placed on a carpet is a mild NSE and breaking a vase on the hard surface is a severe NSE. $\langle v,c\rangle$ are used for training. All instances have unavoidable NSEs.
@@
Push: In this safety-gymnasium domain, the robot aims to push a box quickly to a goal state [37]. Pushing a box on a hazard zone (blue circles) produces NSEs. We modify the domain such that in addition to the existing actions, the agent can also wrap the box that costs $+1$. The NSEs can be avoided by pushing a wrapped box. A state is represented as $\langle x,y,b,w,h\rangle$ where, $x,y$ are the robot’s coordinates, $b$ indicates carrying a box, $w$ indicates if box is wrapped and $h$ denotes if it is a hazard area. $\langle b,w,h\rangle$ are used for training.
Atari Freeway In this Atari game, the robot (a chicken) navigates ten cars moving at varying speeds to reach the destination quickly while avoiding being hit. Being hit by a car moves the robot back to its previous position, and is a severe NSE. A game state is defined by coordinates $(x_{1},y_{1})$ and $(x_{2},y_{2})$, i.e., the top left and bottom right corners of the robot and cars, extracted from the Atari-HEAD dataset [36]. Similar to [10], only car coordinates within a specific range of the robot are considered. The robot can move up, down or stay in place, with unit cost and deterministic transitions.
@@ V-A Results and Discussion
(a) Navigation: Unavoidable NSE
(b) Vase: Unavoidable NSE
(c) Safety-gym Push
(d) Navigation: Unavoidable NSE
(e) Vase: Unavoidable NSE
(f) Safety-gym Push
(g) Navigation: Unavoidable NSE
(h) Vase: Unavoidable NSE
(i) Safety-gym Push
(j) Navigation: Unavoidable NSE
(k) Vase: Unavoidable NSE
(l) Safety-gym Push
Figure 6: Average penalty incurred, along with standard error, when learning from a single feedback (a-c), and using combinations of two formats (d-l).

### v2

@@ IV-A Critical States Selection
Since $p^{t}$ and $q^{t}$ contain categorical values rather than probabilities, their corresponding empirical probability mass functions (PMFs) are computed over the three NSE categories (no NSE, mild NSE, and severe NSE), yielding $\hat{p}^{t}$ and $\hat{q}^{t}$, respectively. In this case, $\hat{p}^{t}$ and $\hat{q}^{t}$ will be vectors of length three, since we consider three NSE categories.
In order to select critical states for querying, we compute the KL divergence between $\hat{q}^{t-1}$ and $\hat{p}^{t}$, $D_{KL}(\hat{p}^{t}\|\hat{q}^{t-1})$. Although $D_{KL}(\hat{p}^{t}\|\hat{q}^{t})$ may appear as a reasonable criterion to guide critical states selection, it only measures how well the agent learns from the feedback at $t$. It does not reveal states where the agent’s predictions were incorrect. For the example shown in Figure 3 with $q^{t-1}(s)=[0,0,0,0]$ and $p^{t}(s)=[2,0,0,0]$, $\hat{p}^{t}$ and $\hat{q}^{t-1}$ are calculated as the average occurrence of each NSE category (no NSE, mild NSE, severe NSE) across the four actions. That is, for $q^{t-1}(s)=[0,0,0,0]$, the frequency is $\frac{4}{4},\frac{0}{4},\frac{0}{4}]$, resulting in $\hat{q}^{t-1}(s)=[1.0,0.0,0.0]$. For $p^{t}(s)=[2,0,0,0]$, the frequency is $\frac{3}{4},\frac{0}{4},\frac{1}{4}]$, yielding $\hat{p}^{t}(s)=[0.75,0.0,0.25]$. Calculating the divergence between $\hat{p}^{t}(s)$ and $\hat{q}^{t-1}(s)$ reveals that the prediction was incorrect at $s$ and therefore more data is required to align the learned model, and hence $s$ or similar states should be selected for querying. Therefore, the sampling weight of the cluster containing $s$ is increased (the region where the NSE model is still uncertain). In the following iteration, critical states are drawn from the reweighted clusters. Algorithm 1 outlines our approach for selecting critical states at each learning iteration, with the following three key steps.
2. Estimating information gain: We define the information gain of sampling from a cluster $k\!\in\!K$, based on the learning gap, as follows:
$IG(k)^{t}\!$ $=\!\frac{1}{|\Omega_{k}^{t-1}|}\sum_{s\in\Omega_{k}^{t-1}}D_{KL}(\hat{p}^{t}\|\hat{q}^{t-1})$ (1)
$=\!\frac{1}{|\Omega_{k}^{t-1}|}\sum_{s\in\Omega_{k}^{t-1}}\sum_{l\in\{l_{a},l_{m},l_{h}\}}\hat{p}^{t}(l|s)\cdot\log\left(\frac{\hat{p}^{t}(l|s)}{\hat{q}^{t-1}(l|s)}\right),$ (2)
where $\Omega_{k}^{t-1}$ denotes the set of states sampled for querying from cluster $k$ at iteration t1t-1. $\hat{p}^{t}(l|s)$ and $\hat{q}^{t-1}(l|s)$ denote the probability of observing NSE category $l\in\{l_{a},l_{m},l_{h}\}$ in state $s$, derived from $p^{t}$ and $q^{t}$, respectively. This formulation quantifies how much the predicted NSE distribution diverges from the feedback received for each state, providing a principled measure of the expected information gain from querying in a cluster, $k$.
3. Sampling critical states: At each learning iteration $t$, the agent assigns a weight $w_{k}$ to each cluster $k\!\in\!K$, proportional to the new information on NSEs revealed by the most informative feedback format identified at t1t-1, using Eqn. 2. Clusters are given equal weights when there is no prior feedback (Line 4). Let $N$ denote the number of critical states to be sampled in every iteration. We sample critical states in batches but they can also be sampled sequentially. When sampling in batches of $N$ states, the number of states $n_{k}$ to be sampled from each cluster is determined by its assigned weight. At least one state is sampled from each cluster to ensure sufficient information for calculating the information gain for every cluster (Line 5). The agent randomly samples $n_{k}$ states from corresponding cluster and adds them to a set of critical states $\Omega$ (Lines 6, 7). If the total number of critical states sampled is less than $N$ due to rounding, then the remaining $N_{r}$ states are sampled from the cluster with the highest weight and added to $\Omega$ (Lines 9-12).
@@
0: B, Querying budget; $D$, Human preference model; $\delta$: KL divergence threshold
$1t\leftarrow$ 1; $\mathcal{V}_{f}\leftarrow 0$ and $n_{f}\leftarrow 0,\ \forall f\in\mathcal{F}$
[…]
3:while $B>0$ or $\forall s\in S,D_{KL}(\hat{p}^{t}\|\hat{q}^{t})\leq\delta$ do
@@
[…]
9:   $q^{t}\leftarrow\{\mathcal{P}(s,a),\forall a\in A,\forall s\in\Omega\}$
@@
IV-B Feedback Format Selection
To query in the critical states, $\Omega$, it is important to select a feedback format that not only maximizes the expected information gain about NSEs but also accounts for likelihood and cost of the feedback. The information gain of a feedback format $f$ at iteration $t$, for $N\!=\!|\Omega|$ critical states, is computed as the KL divergence between the observed and predicted NSE severity distributions, $\hat{p}^{t}$ and $\hat{q}^{t}$:
$\mathcal{V}_{f}$ $=\frac{1}{N}\sum_{s\in\Omega}D_{KL}(\hat{p}^{t}\|\hat{q}^{t})\cdot\mathbb{I}[f\!=\!f_{H}^{t}]+\mathcal{V}_{f}\cdot(1-\mathbb{I}[f\!=\!f_{H}^{t}]),$ (3)
where, $\mathbb{I}[f\!=\!f_{H}^{t}]$ is an indicator function that checks whether the format provided by the human, $f_{H}^{t}$, matches the requested format $f$. If no feedback is received, the information gain for that format remains unchanged. The following equation is used to select the feedback format $f^{*}$, accounting for feedback cost and user preferences:
[…]
where $\psi(f)$ is the probability of receiving a feedback in format $f$ and $C(f)$ is the feedback cost, determined using the human preference model $D$. $t$ is the current learning iteration, $n_{f}$ is the number of times feedback in format $f$ was received, and $\epsilon$ is a small constant for numeric stability. The selected format $f^{*}$ represents the most informative feedback format given the agent’s current knowledge, balancing exploration (less frequently used formats) and exploitation (formats known to provide high information gain).
@@ IV-B Feedback Format Selection
Figure 5: Illustrations of evaluation domains. Red box denotes the agent and the goal location is in green.
Algorithm 2 outlines our feedback format selection approach. Since the agent has no prior knowledge of how the human categorizes NSE for each state-action pairs, the labeling function $p$ is instantiated by sampling from a uniform prior over the three NSE labels ($l_{a},l_{m},l_{h}$) for every $(s,a)$, while q is initialized assuming all actions are safe ($l_{a}$) (Line 2). These initial labels are progressively refined as human feedback is received. At each iteration, the agent samples $|\Omega|$ critical states using Algorithm 1 (Line 4), and selects a feedback format $f^{*}$ is selected using Eqn. 4. The agent queries the human for feedback in $f^{*}$ (Line 5). If the feedback is received (with probability $\psi(f^{*})$), the observed NSE labels $p^{t}$ are updated and an NSE prediction model $\mathcal{P}$ is trained (Lines 6-8). The classifier $\mathcal{P}$ predicts the labels for the sampled critical states $\Omega$, yielding $q^{t}$. We restrict the prediction to $\Omega$ since these states indicate regions of high uncertainty and contribute to reducing the divergence between the true and learned NSE distributions. Further, restricting predictions to $\Omega$ also reduces computational overhead during iterative querying. $\mathcal{V}_{f^{*}}$ recomputed using Eqn. 3, and $n_{f^{*}}$ is incremented (Lines 9-11). This repeats until either the querying budget is exhausted or the KL divergence between $\hat{p}^{t}$ and $\hat{q}^{t}$ over all states is within a problem-specific threshold $\delta$.
Figure 6 illustrates the critical states and the most informative feedback formats selected at each iteration in the object delivery task using AFS, demonstrating that feedback utility changes over time, based on the robot’s current knowledge.
IV-C Stopping Criteria
Besides guiding the selection of critical states and feedback format, the KL-divergence also serves as an indicator of when to stop querying. The querying phase can be terminated when $D_{KL}(\hat{p}^{t}\|\hat{q}^{t})\leq\delta$, where $\delta$ is a problem-specific threshold. When $D_{KL}(\hat{p}^{t}\|\hat{q}^{t})\leq\delta$, it indicates that the learned model is a reasonable approximation of the underlying NSE distribution, and therefore the querying can be terminated even if the allotted budget $B$ has not been exhausted. The choice of $\delta$ provides a trade-off between thorough learning and human effort, and can be tuned based on domain-specific safety requirements.
Figure 6: Feedback utility of each format across iterations. Numbers mark when a state was identified as critical, and circle colors denote the chosen feedback format.
V EXPERIMENTS IN SIMULATION
We first evaluate AFS on three simulated domains (Fig. 5). Human feedback is simulated by modeling an oracle that selects safer actions with higher probability using a softmax action selection [16, 22]: the probability of choosing an action $a^{\prime}$ from a set of all safe actions $A^{*}$ in state $s$ is, $\Pr(a^{\prime}|s)=\frac{e^{Q(s,a^{\prime})}}{\sum\limits_{a\in A^{*}}e^{Q(s,a)}}$.
@@ V EXPERIMENTS IN SIMULATION
Baselines (i) Naive Agent: The agent naively executes its primary policy without learning about NSEs, providing an upper bound on the NSE penalty incurred. (ii) Oracle: The agent has complete knowledge about $R_{T}$ and $R_{N}$, providing a lower bound on the NSE penalty incurred. (iii) Reward Inference with $\beta$ Modeling (RI) [16]: The agent selects a feedback format that maximizes information gain according to the human’s inferred rationality, $\beta$. (iv) Cost-Sensitive Approach: The agent selects a feedback method with the least cost, according to the preference model $D$. (v) Most-Probable Feedback: The agent selects a feedback format that the human is most likely to provide, based on $D$. (vi) Random Critical States: The agent uses our AFS framework to learn about NSEs, except the critical states are sampled randomly from the entire state space. We use $\theta_{1}\!=\!1$ and $\theta_{2}\!=\!1$ for all our experiments. AFS uses learned $\hat{R}_{N}$.
@@
Domains, Metrics and Feedback Formats We evaluate the performance of various techniques on three domains in simulation (Figure  5): outdoor navigation, vase and safety-gym’s push. We optimize costs (negations of rewards) and compare techniques using average NSE penalty and average cost to goal, averaged over 100 trials. For navigation, vase and push, we simulate human feedback. The cost for $l_{a}$, $l_{m}$, and $l_{h}$ are 0, +5+5, and +10+10 respectively.
Figure 7: Average penalty incurred when learning with AFS using querying budget $B=400$, and KL divergence (KLD) as the stopping criterion. The budget utilized by AFS with KLD stopping is annotated in the plot.
Figure 8: Average penalty incurred when querying with different feedback selection techniques.
@@
Navigation: In this ROS-based city environment, the robot optimizes the shortest path to the goal location. A state is represented as $\langle x,y,f,p\rangle$, where, $x$ and $y$ are robot coordinates, $f$ is the surface type (concrete or grass), and $p$ indicates the presence of a puddle. The robot can move in all four directions and each costs +1+1. Actions succeed with probability $0.8$. Navigating over grass damages the grass and is a mild NSE. Navigating over grass with puddles is a severe NSE. Features used for training are $\langle f,p\rangle$. Here, NSEs are unavoidable.
@@
Vase: In this domain, the robot must quickly reach the goal, while minimizing breaking a vase as a side effect [27]. A state is represented as $\langle x,y,v,c\rangle$ where, $x$ and $y$ are robot’s coordinates. $v$ indicates the presence of a vase and $c$ indicates if the floor is carpeted. The robot moves in all four directions and each costs +1+1. Actions succeed with probability $0.8$. Breaking a vase placed on a carpet is a mild NSE and breaking a vase on the hard surface is a severe NSE. $\langle v,c\rangle$ are used for training. All instances have unavoidable NSEs.
@@
Push: In this safety-gymnasium domain, the robot aims to push a box quickly to a goal state [23]. Pushing a box on a hazard zone (blue circles) produces NSEs. We modify the domain such that in addition to the existing actions, the agent can also wrap the box that costs +1+1. Every move action succeeds with probability 0.8, and the wrap action succeeds with probability 1.0. The NSEs can be avoided by pushing a wrapped box. A state is represented as $\langle x,y,b,w,h\rangle$ where, $x,y$ are the robot’s coordinates, $b$ indicates carrying a box, $w$ indicates if box is wrapped and $h$ denotes if it is a hazard area. $\langle b,w,h\rangle$ are used for training.
@@ V-A Results and Discussion
Effect of learning using AFS: We first examine the benefit of querying using AFS, by comparing the resulting average NSE penalties and the cost for task completion, across domains and query budget. Figure 8 shows the average NSE penalties when operating based on an NSE model learned using different querying approaches. Clusters for critical state selection were generated using KMeans clustering algorithm with K=3K\!=\!3 for navigation, vase and safety-gym’s push domains (Figure 8 (a-c)). The results show that our approach consistently performs similar to or better than the baselines.
There is a trade-off between optimizing task completion and mitigating NSEs, especially when NSEs are unavoidable. While some techniques are better at mitigating NSEs, they significantly impact task performance. Table I shows the average cost for task completion at B=400B\!=\!400. Lower values are better for both NSEs and task completion cost. While the Naive Agent has a lower cost for task completion, it incurs the highest NSE penalty as it has no knowledge of $R_{N}$. RI causes more NSEs, especially when they are unavoidable, as its reward function does not fully model the penalties for mild and severe NSEs. Overall, the results show that our approach consistently mitigates avoidable and unavoidable NSEs, without affecting the task performance substantially.
Figure 7 shows the average penalty when AFS uses KL-divergence (KLD) as the stopping criteria, compared to querying with budget $B=400$. For comparison, we also annotate in the plot the querying budget used by AFS with KLD stopping at the time of termination. The results show that despite terminating earlier and using few queries, AFS with the KLD stopping achieves comparable performance to that of AFS with query budget $B=400$, demonstrating the usefulness of KLD as a stopping criterion.
Method Navigation: unavoidable NSE Vase: unavoidable NSE Safety-gym Push: avoidable NSE
Oracle $51.37\pm 2.69$ $54.46\pm 6.70$ $44.62\pm 9.97$
Naive $36.11\pm 1.39$ $36.0\pm 2.89$ $39.82\pm 5.44$
RI $40.10\pm 0.69$ $37.42\pm 1.01$ $42.15\pm 2.44$
AFS $64.8\pm 2.3$ $52.68\pm 7.87$ $48.32\pm 4.42$